Bacteria grow by cell division. If each cell divides into 2 in every 2 minutes how many cells will exist after 32 minutes? Assume there was only one cell at the beginning
Answer by nerdybill(7384)

Bacteria grow by cell division. If each cell divides into 2 in every 2 minutes how many cells will exist after 32 minutes? Assume there was only one cell at the beginning
.
Apply exponential growth equation:
A(t) = Pe^(rt)
where
A(t) is amount at time t (2)
P is the initial amount (1)
r is rate of growth (unknown)
t is time (2)
.
Finding r:
A(t) = Pe^(rt)
2 = 1e^(2r)
ln(2) = 2r
ln(2)/2 = r
.
Cells at 32 minutes:
A(32) = 1e^(ln(2)/2 * 32)
A(32) = e^(ln(2) * 16)
A(32) = 65536 cells